Session: Feb 2015 (session 1)
Wavelength: 6 cm
To give feedback on this experiment, click
(BoXia at Sat Mar 07 10:24:04 2015) Success: LATE
Because of the time conflict with CE-5T1 task, this observation started
late at UT 22:25
(Ruben Fernández at Sat Mar 07 10:36:58 2015) Minor failures:
Scans nº344,345 and 351 were not recorded due to problems at the
Tsys high due to problems with the band C compressor.
(Stepanova, Gusev at Sat Mar 07 15:34:07 2015) Success:
(Lysenkova M. at Sat Mar 07 15:34:25 2015) Success:
(Panova, Nekrasov, Alakova at Sat Mar 07 15:34:43 2015) Success:
(Pawel Wolak at Sat Mar 07 15:44:02 2015) Success:
No known problems.
(Hua Zhang at Sun Mar 08 02:22:04 2015) Failure:
Ur did not observed because of Receiver fault.
(Geert Kuper at Sun Mar 08 22:09:27 2015) Success:
(Pablo de Vicente at Mon Mar 09 07:53:48 2015) Failure:RECE
C band receiver got warm due to an error in the cryostat.
(Andrea Orlati at Tue Mar 10 09:01:55 2015) Success:
no known problems
(Uwe at Wed Mar 11 10:01:11 2015) Success:
(Jun Yang at Fri Mar 13 11:11:23 2015) Success:
Local GPS antenna was broken during the observation and fixed in the
morning of the 2nd day.